OpenStudy (anonymous):

Pls help A. 1440 B. 3992 C. 4008 D. 100,000

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

the link

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

somone

OpenStudy (anonymous):

1440 because cross out 12000 with 25 then take 480 times 3 so you get 1440

OpenStudy (dayakar):

POPULATION OF JUNCTION CITY = 3/25 * JAMESBURG POPULATION \[= \frac{ 3 }{ 25 }\times 12000\]

OpenStudy (anonymous):

so is it really 1440?

OpenStudy (dayakar):

VERY GOOD

OpenStudy (anonymous):

ok thx

OpenStudy (dayakar):

:)